Determine the density of oxygen at 156,000 Pa pressure and Temperature = -8 °C. State your answer to nearest 0.001 kg/m3, follow the standard rounding rules in stating answer, and do not attempt to write the units. Take the molar mass of molecular oxygen to be 32 g/mol, R=8.31 JK-1mol-1 and take the absolute zero to be -273 °C.
Answer #1

The density of oxygen is calculated, from the equation of ideal gases:

D = m / v = P * MM / R * T = (156000/101325) atm * 32 g / mol / 0.082 atm * L / mol * K * (-8 + 273) K = 2.267 g / L

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.